A chartered bank has $1 million in deposits and $40,000 in desired reserves. Its excess reserves are initially zero.

a. The reserve ratio in the banking system is .......%.

b. If a further $100,000 is deposited in this bank then the bank's desired reserves increase by $.......while the bank's excess reserves increase by $........

c. The banking system can increase the money supply by this bank's initial amount of $........multiplied by the money multiplier of ...........for a final increase in the money supply of $............
Answer #1

(a)

Deposits = $1 million

Desired reserves = $40,000

Calculate the reserve ratio -

Reserve ratio = Desired reserves/Deposits

Reserve ratio = $40,000/$1,000,000

Reserve ratio = 0.04 or 4%

So,

The reserve ratio in the banking system is 4%.

(b)

New deposit = $100,000

Reserve ratio = 0.04

Desired reserves = New deposits * Reserve ratio = $100,000 * 0.04 = $4,000

Excess reserves = New deposit - Desired reserves = $100,000 - $4,000 = $96,000

So,

If a further $100,000 is deposited in this bank then the bank's desired reserves increases by $4,000 while the bank's excess reserves increases by $96,000.

(c)

Calculate the money multiplier -

Money multiplier = 1/Reserve ratio = 1/0.04 = 25

The money multiplier is 25.

Calculate the final increase in the money supply -

Final increase in money supply = Increase in excess reserves * money multiplier

Final increase in money supply = $96,000 * 25 = $2,400,000

The final increase in the money supply is $2,400,000 or $2.4 million.

Thus,

The banking system can increase the money supply by this bank's initial amount of $96,000 multiplied by the money multiplier of 25 for a final increase in the money supply of $2.4 million.